OpenStudy (anonymous):

Find the slope of the graph of the relation 2x^3 − 3xy + y^3 = −1 at the point (2, −3) negative eleven sevenths 5 over 7 negative 1 over 3 4 over 3

OpenStudy (anonymous):

I think you take the derivative first which is what i believe to be 6x^2-3y=0 but then idk what to do after that. Help Please.

myininaya (myininaya):

well it looks like you didn't use product rule and chain rule correctly.

myininaya (myininaya):

\[\frac{d}{dx}(xy)= y \frac{dx}{dx}+x \frac{dy}{dx} \text{ by product rule } \\ \frac{d}{dx}(xy)=y(1)+xy' \\ c \frac{d}{dx}(xy)=c(y+xy')\]

myininaya (myininaya):

that last rule down there is constant multiple rule apply to the product rule that I used

myininaya (myininaya):

\[\frac{d}{dx}(y^n)=ny^{n-1}y' \text{ by product rule and chain rule }\]

myininaya (myininaya):

@Krissy3039 let me know if you don't understand

OpenStudy (anonymous):

Yeah I still dont understand completely.

myininaya (myininaya):

Which part ?

myininaya (myininaya):

oops I said product rule and chain rule meant power rule and chain rule*

myininaya (myininaya):

Is that the part you don't understand?